# Neural Graph Pattern Machine

### v1

@@ Abstract
Graph learning tasks require models to comprehend essential substructure patterns relevant to downstream tasks, such as triadic closures in social networks and benzene rings in molecular graphs. Due to the non-Euclidean nature of graphs, existing graph neural networks (GNNs) rely on message passing to iteratively aggregate information from local neighborhoods. Despite their empirical success, message passing struggles to identify fundamental substructures, such as triangles, limiting its expressiveness. To overcome this limitation, we propose the Neural Graph Pattern Machine (GPM), a framework designed to learn directly from graph patterns. GPM efficiently extracts and encodes substructures while identifying the most relevant ones for downstream tasks. We also demonstrate that GPM offers superior expressivity and improved long-range information modeling compared to message passing. Empirical evaluations on node classification, link prediction, graph classification, and regression show the superiority of GPM over state-of-the-art baselines. Further analysis reveals its desirable out-of-distribution robustness, scalability, and interpretability. We consider GPM to be a step toward going beyond message passing.
@@ 1 Introduction
Graphs are essential modeling tools for a wide range of applications, such as social network analysis, drug discovery, and recommendation systems. Many real-world tasks can be transformed into graph classification or regression problems. For instance, determining the chemical properties of molecules can be framed as classifying molecule graphs, where atoms are nodes and bonds are edges (Hu et al., 2020). In these graph-related tasks, the graph topology itself often provides critical insights for downstream tasks (Xu et al., 2019; Zhao et al., 2022). For example, triadic closure [^1] This pattern is prevalent in various networks, such as protein-protein interaction networks, e-commerce networks, communication networks, collaboration networks, and web networks. (Granovetter, 1973), where three nodes are interconnected, illustrates that a friend of a friend is also a friend in social networks, highlighting stable relationships among nodes. It is a fundamental pattern for predicting the occupations or interests of individuals (node classification) (Jin et al., 2020) and for recommending new friends (link prediction) (Huang et al., 2015). Another example is the benzene ring, a cyclic hydrocarbon with six carbon atoms, in molecules (Rong et al., 2020). The structure of the benzene ring indicates molecular stability due to delocalized electrons, making it resistant to certain chemical reactions like addition reactions. In this paper, we refer to these substructures (triangles, rings, etc.) as substructure patterns or graph patterns.
@@
To model graph-structured data, graph neural networks (GNNs) (Kipf & Welling, 2017; Hamilton et al., 2017; Veličković et al., 2018) have been introduced and have achieved remarkable success in tasks such as node classification, link prediction, and graph classification. Despite these empirical successes, basic GNNs rely on message passing (Gilmer et al., 2017) to iteratively aggregate neighborhood information and update node representations, rather than directly learning from graph patterns. Therefore, numerous studies (Xu et al., 2019; Verma & Zhang, 2019; Garg et al., 2020; Chen et al., 2020; Tang & Liu, 2023; Zhang et al., 2024a) have shown that message passing struggles to identify even the most fundamental patterns, such as triangles, stars, and $k$-cliques, due to its limitation to the 1-dimensional Weisfeiler-Leman (1-WL) isomorphism test. To enhance the pattern modeling capabilities of message passing, advanced methods such as positional embeddings (Murphy et al., 2019; Loukas, 2020), graph transformers (Kreuzer et al., 2021; Rampasek et al., 2022), and expressive GNNs (Morris et al., 2019; Zhao et al., 2022) have been proposed, breaking the 1-WL test bound. However, these methods essentially complement message passing and cannot yet directly learn from graph patterns. Additionally, they encounter other limitations, such as lack of interpretability, quadratic complexity, and biased inductive biases.
@@
To go beyond the existing message-passing frameworks, researchers have begun exploring the use of graph patterns as tokens to represent graph instances (nodes, links, graphs). However, this area of research is still in its infant stage. There are three main challenges in designing such methods. (1) Universal Graph Tokenizer: Extracting graph patterns requires a tokenizer that can adapt to node-, link-, and graph-level tasks. Unlike sentences and images, which inherently have sequential structures, tokenizing graph instances is challenging due to their non-Euclidean nature. Existing tokenizers tend to be task-specific, such as the Hop2token tokenizer (Chen et al., 2023) for node tasks and the METIS tokenizer (He et al., 2023) for graph tasks. (2) Effective Pattern Encoder: The model must efficiently and comprehensively encode the extracted graph patterns. Current approaches (Chen et al., 2022; He et al., 2023; Behrouz & Hashemi, 2024) often use message passing as a pattern encoder or auxiliary module (Behrouz & Hashemi, 2024) to provide graph inductive biases. However, the limited expressiveness of message passing in identifying basic substructures leads to information loss in encoding graph patterns. (3) Important Pattern Identifier: As the sampled patterns may be duplicated or noisy, it is crucial to identify the most relevant patterns for downstream domains. For example, in social networks that favor localized patterns, the model should emphasize patterns that preserve local information. However, existing methods are generally evaluated on benchmarks with long-range dependencies (Dwivedi et al., 2023, 2022; Platonov et al., 2023); their effectiveness on graphs favoring localized or mixed dependencies remains unclear.
@@
To address these challenges, we introduce the Neural Graph Pattern Machine (GPM), where the workflow is illustrated in Figure 1. We design a random walk-based tokenizer to sample graph patterns, which is computationally efficient (Grover & Leskovec, 2016) and can be naturally adapted to various tasks. The key insight is that “the combination of semantic path and anonymous path of a random walk matches a particular graph pattern” (proved in Section 3.1). By leveraging this insight, we model graph patterns by separately encoding the semantic paths and anonymous paths of the corresponding walks, thereby comprehensively capturing the preserved graph inductive biases. The encoded graph patterns are then fed into a transformer layer (Vaswani et al., 2017) that identifies the important patterns dominating downstream tasks. To provide a deeper understanding on the superiority of GPM over message passing, we demonstrate that GPM can distinguish non-isomorphic graphs that GNNs cannot identify and can model long-range dependencies that GNNs fail to capture. Furthermore, we conduct extensive experiments on node-, link-, and graph-level tasks to demonstrate that GPM is applicable to various graph tasks and outperforms state-of-the-art baselines. Our experimental results also show that GPM is robust to out-of-distribution issues, and can be readily scaled to large graphs, large model sizes, and distributed training. In addition, the mechanism for identifying dominant substructures enables GPM to have desirable model interpretability.
@@ 2 Related Works
We primarily discuss methods that utilize graph patterns as tokens in this section. A detailed discussion of the limitations of message passing and existing improvements is provided in Appendix A. Early methods for leveraging graph patterns often relied on primitive training paradigms. For instance, DGK (Yanardag & Vishwanathan, 2015) uses graph kernels to measure relationships between (substructure) patterns, while AWE (Ivanov & Burnaev, 2018) employs anonymous random walks to encode pattern distributions in graphs. More recent approaches tokenize graphs into sequences of substructures. GraphViT (He et al., 2023), for example, splits a graph into multiple subgraphs using graph partitioning algorithms (Karypis, 1997), encodes each subgraph individually via message passing, and aggregates the resulting embeddings to represent the entire graph. Similarly, GMT (Baek et al., 2021) decomposes nodes into a multiset, with each set representing a specific graph pattern. However, these methods are limited to graph-level tasks. On the other hand, NAGphormer (Chen et al., 2023) and GCFormer (Chen et al., 2024) utilize Hop2token and neighborhood sampling, respectively, to tokenize patterns representing individual nodes; Yet, they are tailored for node-level tasks. To address this, SAT (Chen et al., 2022), GNN-AK (Zhao et al., 2022), and GraphMamba (Behrouz & Hashemi, 2024) propose task-agnostic tokenization methods. However, these methods often rely on message passing either as the pattern encoder or as an auxiliary module, inheriting the limitations of message passing. While the aforementioned methods have shown empirical success, none fully meet our criteria: (1) a universal graph tokenizer, (2) an effective pattern encoder, and (3) the ability to identify important patterns relevant to downstream tasks. Our proposed GPM overcomes these challenges.
@@ 3 Neural Graph Pattern Machine
Let ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$ denote a graph, where ${\mathcal{V}}$ is the set of nodes with $|{\mathcal{V}}|=N$, and ${\mathcal{E}}\subseteq{\mathcal{V}}\times{\mathcal{V}}$ is the set of edges with $|{\mathcal{E}}|=E$. Each node $v\in{\mathcal{V}}$ is associated with a feature vector $\mathbf{x}\in\mathbb{R}^{d_{n}}$, and each edge $e\in{\mathcal{E}}$ (if applicable) is associated with a feature vector $\mathbf{e}\in\mathbb{R}^{d_{e}}$. We define graph patterns as subgraphs that represent small, recurring substructures within graphs, such as triangles, stars, cycles, etc.
@@ 3.1 Pattern Tokenizer
Given a node $v\in{\mathcal{V}}$, assume the task requires information from the $k$-hop ego-graph ${\mathcal{B}}(v,k)=({\mathcal{V}}^{\prime},{\mathcal{E}}^{\prime})$. A sufficiently large set of patterns, sampled via $l$-length anonymous walks with $l=O(|{\mathcal{E}}^{\prime}|)$, can provide distinguishable topological representations.
The proof is sketched in Appendix C.1. This proposition demonstrates that the distribution of anonymous paths starting from a node is sufficient to capture its topological properties. Naturally, two nodes $u$ and $v$ can be considered to share similar graph patterns if their anonymous path distributions are similar. This proposition can be extended to links and graphs by treating them as combinations of nodes, with the corresponding distributions. Thus, for each random walk, we can derive both a semantic path (capturing node features) and an anonymous path (capturing topological structures), where their combination conceptually represents a certain graph pattern. Then, we have the following proposition.
@@
Jointly encoding semantic paths and anonymous paths via any bijective mappings provides a comprehensive representation of the graph inductive biases encapsulated in the corresponding graph pattern.
@@ 3.2 Pattern Encoder
where ${\bm{x}}_{i}$ denotes the node features, ${\bm{e}}_{i-1,i}$ represents optional edge features, and $\alpha_{i}$ refers to the optional positional embedding. In this work, we utilize widely adopted PEs, including random-walk structural embeddings (RWSE) (Rampasek et al., 2022) and Laplacian eigenvector embeddings (Lap) (Kreuzer et al., 2021). Empirically, GPM achieves competitive performance without the use of PEs.
@@ 3.3 Important Pattern Identifier
where ${\bm{W}}_{{\bm{Q}}},{\bm{W}}_{{\bm{K}}},{\bm{W}}_{{\bm{V}}}$ are trainable parameter matrices, $d_{out}$ is the dimension of the query matrix ${\bm{Q}}$, and $\operatorname{FFN}$ represents a two-layer MLP. We utilize multi-head attention, which has proven effective in practice by concatenating multiple attention mechanisms.
@@ 3.5 How Does GPM Surpass Message Passing?
GPM can distinguish non-isomorphic graphs given a sufficient number of graph patterns.
The proof is sketched by first demonstrating the expressive power of a simplified variant of GPM and then generalizing to the full model. The detailed proof is in Appendix C.3.
Theorem 3.5.
@@
The proof of this result is detailed in Appendix C.4. This theorem highlights that GPM is at least as expressive as message passing frameworks and high-order GNNs. Empirically, GPM successfully distinguishes graphs that remain indistinguishable under the 3-WL test (Appendix F.1).
Tackling Over-Squashing. Another limitation of message passing is their focus on localized information, which prevents them from effectively capturing long-range dependencies within graphs. In contrast, GPM demonstrates superior capability in modeling long-range interactions. Following He et al. (2023), we evaluate this capability using the TreeNeighborsMatching (Alon & Yahav, 2021). Our GPM perfectly fits the data with task radius up to 7, yet message passing methods exhibit over-squashing effects as early as task radius 4 (Appendix F.2).

### v2

# Beyond Message Passing: Neural Graph Pattern Machine

@@ Abstract
Graph learning tasks often hinge on identifying key substructure patterns—such as triadic closures in social networks or benzene rings in molecular graphs—that underpin downstream performance. However, most existing graph neural networks (GNNs) rely on message passing, which aggregates local neighborhood information iteratively and struggles to explicitly capture such fundamental motifs, like triangles, $k$-cliques, and rings. This limitation hinders both expressiveness and long-range dependency modeling. In this paper, we introduce the Neural Graph Pattern Machine (GPM), a novel framework that bypasses message passing by learning directly from graph substructures. GPM efficiently extracts, encodes, and prioritizes task-relevant graph patterns, offering greater expressivity and improved ability to capture long-range dependencies. Empirical evaluations across four standard tasks—node classification, link prediction, graph classification, and graph regression—demonstrate that GPM outperforms state-of-the-art baselines. Further analysis reveals that GPM exhibits strong out-of-distribution generalization, desirable scalability, and enhanced interpretability. Code and datasets are available at: https://github.com/Zehong-Wang/GPM.
@@ 1 Introduction
Graphs serve as a fundamental abstraction for modeling complex systems across domains such as social networks, drug discovery, and recommender systems (Zhang et al., 2024c, d; Ma et al., 2025a; Zhang et al., 2024b; Ma et al., 2023; Fu et al., 2023). Many real-world problems can be formulated as classification or regression tasks on graphs. For example, molecular property prediction can be cast as a graph classification task, where atoms and bonds are represented as nodes and edges, respectively (Hu et al., 2020). In e-commerce, predicting user preferences naturally translates into link prediction on user-item interaction graphs (Ying et al., 2018). A key insight of these tasks is that the certain substructures encode meaningful inductive biases (Xu et al., 2019; Zhao et al., 2022), emerging as predictive patterns. For instance, triadic closure, where three nodes form a closed triangle, is ubiquitous in social, biological, and communication networks (Granovetter, 1973). It serves as an indicator of stable node relationships and is instrumental in tasks like node classification (Jin et al., 2020) and link prediction (Huang et al., 2015). In molecular graphs, the benzene ring—a six-carbon cyclic structure—is a canonical example of a stable substructure with implications for chemical reactivity (Rong et al., 2020). We refer to such recurring substructures, including triangles, rings, and other motifs, as substructure patterns or graph patterns. These patterns form the building blocks of graph semantics and are central to understanding and improving performance in graph-based learning tasks.
@@
Despite the critical role of substructure patterns in graph learning, most graph neural networks (GNNs) (Kipf & Welling, 2017; Hamilton et al., 2017; Veličković et al., 2018) operate under the message passing paradigm (Gilmer et al., 2017), which iteratively aggregates local neighborhood information rather than directly modeling graph substructures. While message passing GNNs have demonstrated strong empirical performance on tasks such as node classification, link prediction, and graph classification, numerous studies (Xu et al., 2019; Verma & Zhang, 2019; Garg et al., 2020; Chen et al., 2020; Tang & Liu, 2023; Zhang et al., 2024a) highlight their inherent limitations in capturing basic patterns like triangles, stars, and $k$-cliques, owing to their equivalence to the 1-dimensional Weisfeiler-Leman (1-WL) test (Xu et al., 2019). To address these limitations, various enhancements have been proposed, including positional encodings (Murphy et al., 2019; Loukas, 2020), graph transformers (Kreuzer et al., 2021; Rampasek et al., 2022), and higher-order GNNs (Morris et al., 2019; Zhao et al., 2022; Qian et al., 2024; Ma et al., 2025b), which extend expressiveness beyond 1-WL. However, these methods remain fundamentally tied to message passing, and often suffer from challenges such as limited interpretability, high computational cost, and biased inductive assumptions.
@@
To go beyond the message passing, recent research has begun exploring the use of graph patterns as discrete tokens to represent nodes, links, or entire graphs. While promising, this line of work is still in its early stages and faces three key challenges. (1) Universal Graph Tokenizer: Extracting graph patterns requires a tokenizer that can adapt to node-, link-, and graph-level tasks. Unlike sentences and images, which inherently have sequential structures, tokenizing graph instances is challenging due to their non-Euclidean nature. Existing tokenizers tend to be task-specific, such as the Hop2token tokenizer (Chen et al., 2023) for node tasks and the METIS tokenizer (He et al., 2023) for graph tasks. (2) Effective Pattern Encoder: The model must efficiently and comprehensively encode the extracted graph patterns. Current approaches often use message passing as a pattern encoder (Chen et al., 2022; He et al., 2023; Behrouz & Hashemi, 2024) or auxiliary module (Behrouz & Hashemi, 2024) to provide graph inductive biases. However, the limited expressiveness of message passing in identifying basic substructures leads to information loss in encoding graph patterns. (3) Important Pattern Identifier: As the sampled patterns may be duplicated or noisy, it is crucial to identify the most relevant patterns for downstream domains. For example, in social networks that favor localized patterns, the model should emphasize patterns that preserve local information. However, existing methods are generally evaluated on benchmarks with long-range dependencies (Dwivedi et al., 2023, 2022; Platonov et al., 2023); their effectiveness on graphs favoring localized or mixed dependencies remains unclear.
@@
To address these challenges, we introduce the Neural Graph Pattern Machine (GPM), where the workflow is illustrated in Figure 1. We design a random walk-based tokenizer to sample graph patterns, which is computationally efficient (Grover & Leskovec, 2016) and can be naturally adapted to various tasks. The key insight is that the combination of semantic path and anonymous path of a random walk matches a particular graph pattern (proved in Section 3.1). By leveraging this insight, we model graph patterns by separately encoding the semantic paths and anonymous paths of the corresponding walks, thereby comprehensively capturing the preserved graph inductive biases. The encoded graph patterns are then fed into a transformer layer (Vaswani et al., 2017) that identifies the important patterns dominating downstream tasks. To provide a deeper understanding on the superiority of GPM over message passing, we demonstrate that GPM can distinguish non-isomorphic graphs that GNNs cannot identify and can model long-range dependencies that GNNs fail to capture. Furthermore, we conduct extensive experiments on node-, link-, and graph-level tasks to demonstrate that GPM is applicable to various graph tasks and outperforms state-of-the-art baselines. Our experimental results also show that GPM is robust to out-of-distribution issues, and can be readily scaled to large graphs, large model sizes, and distributed training. In addition, the mechanism for identifying dominant substructures enables GPM to have desirable model interpretability.
@@ 2 Related Works
Expressive Bottleneck of Message Passing. Traditional message passing GNNs (Kipf & Welling, 2017; Hamilton et al., 2017; Veličković et al., 2018) suffer from well-known issues, particularly on limited expressiveness. Their expressiveness is fundamentally bounded by the 1-WL test (Xu et al., 2019; Corso et al., 2020), which restricts the ability to distinguish basic structures such as stars, cycles, and cliques (Chen et al., 2020; Garg et al., 2020; Zhang et al., 2024a; Qian et al., 2025). To address this limitation, three main research directions have emerged. First, expressive GNN variants enhance the message passing framework by incorporating higher-order structure encodings (Maron et al., 2019a; Bouritsas et al., 2022) and node positional embeddings (Murphy et al., 2019; Loukas, 2020), but often suffer from scalability challenges (Azizian & marc lelarge, 2021). Second, random walk-based approaches (Zhang et al., 2019; Fan et al., 2022; Jin et al., 2022; Wang & Cho, 2024) improve long-range modeling (Welke et al., 2023), but often sacrifice local pattern understanding (Tönshoff et al., 2023). Third, graph transformers (Kreuzer et al., 2021; Ying et al., 2021) utilize global attention to capture arbitrary node dependencies and exceed WL expressiveness, but incur quadratic complexity, limiting scalability to large graphs (Wu et al., 2023). While these approaches extend message passing in expressiveness, they are often implemented as extensions or complements to the message passing, failing to directly and effectively model graph patterns. This observation motivates the need for a fundamentally different approach, as pursued in this work. A more detailed discussion is provided in Appendix A.
Graph Patterns as Tokens. Early methods for leveraging graph patterns often relied on primitive training paradigms. For instance, DGK (Yanardag & Vishwanathan, 2015) uses graph kernels to measure relationships between (substructure) patterns, while AWE (Ivanov & Burnaev, 2018) employs anonymous random walks to encode pattern distributions in graphs. More recent approaches tokenize graphs into sequences of substructures. GraphViT (He et al., 2023), for example, splits a graph into multiple subgraphs using graph partitioning algorithms (Karypis, 1997), encodes each subgraph individually via message passing, and aggregates the resulting embeddings to represent the entire graph. Similarly, GMT (Baek et al., 2021) decomposes nodes into a multiset, with each set representing a specific graph pattern. However, these methods are limited to graph-level tasks. On the other hand, NAGphormer (Chen et al., 2023) and GCFormer (Chen et al., 2024) utilize Hop2token and neighborhood sampling, respectively, to tokenize patterns representing individual nodes; Yet, they are tailored for node-level tasks. To address this, SAT (Chen et al., 2022), GNN-AK (Zhao et al., 2022), and GraphMamba (Behrouz & Hashemi, 2024) propose task-agnostic tokenization methods. However, these methods often rely on message passing either as the pattern encoder or as an auxiliary module, inheriting the limitations of message passing. While the aforementioned methods have shown empirical success, none fully meet our criteria: (1) a universal graph tokenizer, (2) an effective pattern encoder, and (3) the ability to identify important patterns relevant to downstream tasks. Our proposed GPM overcomes these challenges.
@@ 3 Neural Graph Pattern Machine
Let ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}},{\bm{X}},{\bm{E}})$ denote a graph, where ${\mathcal{V}}$ is the set of nodes with $|{\mathcal{V}}|=N$, ${\mathcal{E}}\subseteq{\mathcal{V}}\times{\mathcal{V}}$ is the set of edges with $|{\mathcal{E}}|=E$, and ${\bm{X}}$ and ${\bm{E}}$ represent the node and edge feature matrices, respectively. Each node $v\in{\mathcal{V}}$ is associated with a feature vector $\mathbf{x}_{v}\in\mathbb{R}^{d_{n}}$, and each edge $e\in{\mathcal{E}}$ is associated with a feature vector $\mathbf{e}_{e}\in\mathbb{R}^{d_{e}}$ (if applicable). We define graph patterns as subgraphs that represent small, recurring substructures, such as triangles, stars, cycles, etc.
@@ 3.1 Pattern Tokenizer
(Informal) Given a node $v\in{\mathcal{V}}$, assume the task requires information from the $k$-hop ego-graph ${\mathcal{B}}(v,k)=({\mathcal{V}}^{\prime},{\mathcal{E}}^{\prime})$. A sufficiently large set of patterns, sampled via $l$-length anonymous walks with $l=O(|{\mathcal{E}}^{\prime}|)$, can provide distinguishable topological representations.
The proof is sketched in Appendix C.1. This proposition demonstrates that the distribution of anonymous paths starting from a node is sufficient to capture its topological properties. Naturally, two nodes $u$ and $v$ can be considered to share similar graph patterns if their anonymous path distributions are similar. This proposition can be extended to links and graphs by treating them as combinations of nodes, with the corresponding distributions. Thus, for each random walk, we can derive both a semantic path (capturing node features) and an anonymous path (capturing topological structures), where their combination conceptually represents a certain graph pattern. Then, we have the proposition.
@@
(Informal) Jointly encoding semantic paths and anonymous paths via any bijective mappings provides a comprehensive representation of the graph inductive biases encapsulated in the corresponding graph pattern.
@@ 3.2 Pattern Encoder
where ${\bm{x}}_{i}$ denotes the node features, ${\bm{e}}_{i-1,i}$ represents optional edge features, and $\alpha_{i}$ refers to the optional positional embedding. In this work, we utilize widely adopted PEs, including random-walk structural embeddings (RWSE) (Rampasek et al., 2022) and Laplacian eigenvector embeddings (Lap) (Kreuzer et al., 2021). Empirically, we observe that the choice of PE depends on the dataset. Notably, the model can still achieve competitive performance even without the positional embeddings.
@@ 3.3 Important Pattern Identifier
where ${\bm{W}}_{{\bm{Q}}},{\bm{W}}_{{\bm{K}}},{\bm{W}}_{{\bm{V}}}$ are trainable parameter matrices, $d_{out}$ is the dimension of the query matrix ${\bm{Q}}$, and $\operatorname{FFN}$ represents a two-layer MLP. We utilize multi-head attention, which has proven effective in practice by concatenating multiple attention mechanisms. Additionally, multiple transformer layers can be stacked to further enhance the capacity.
@@ 3.5 How Does GPM Surpass Message Passing?
Under the reconstruction conjecture assumption, GPM can distinguish all pairs of non-isomorphic graphs given a sufficient number of graph patterns.
The proof is sketched by first demonstrating the expressive power of a simplified variant of GPM and then generalizing to the full model. The detailed proof is in Appendix C.3. Based on the theorem, it is readily to extend to the following corollary.
Corollary 3.5.
@@
The proof of this result is detailed in Appendix C.4. This corollary highlights that GPM is at least as expressive as message passing frameworks and high-order GNNs. Empirically, GPM successfully distinguishes graphs that remain indistinguishable under the 3-WL test (Appendix E.1).
Tackling Over-Squashing. Another limitation of message passing is their focus on localized information, which prevents them from effectively capturing long-range dependencies within graphs. In contrast, GPM demonstrates superior capability in modeling long-range interactions. Following He et al. (2023), we evaluate this capability using the TreeNeighborsMatching (Alon & Yahav, 2021). Our GPM perfectly fits the data with task radius up to 7, yet message passing methods exhibit over-squashing effects as early as task radius 4 (Appendix E.2).